# DNABERT-2: Efficient Foundation Model and Benchmark For Multi-Species Genome

### 26 Jun 2023 v1

@@ 3.2 Model
DNABERT-2 adapts the Transformer Encoder architecture similar to BERT [Kenton and Toutanova, 2019]. To address the limitations of existing models, we incorporate a series of recent advances in deep learning to increase the model’s efficiency and capability, including: 1) replacing learned positional embeddings with the Attention with Linear Biases (ALiBi) [Press et al., 2021] to overcome the input length limitation; 2) utilizing FlashAttention [Dao et al., 2022] and Low Precision Layer Normalization to increase computation and memory efficiency; 3) employing the Low-Rank Adaptation (LoRA) [Hu et al., 2021] in the fine-tuning stage (if necessary) for parameter-efficient training.
@@
Due to the permutation-invariant nature of the attention mechanism, explicit positional information is required in attention-based models. Existing solutions such as Sinusoidal [Vaswani et al., 2017], learned [Kenton and Toutanova, 2019], and Rotary [Su et al., 2021] positional embedding methods either suffer from input length restriction or poor extrapolation capability when applied to sequences longer than training data. Attention with Linear Biases (ALiBi) provides an efficient yet effective solution. Instead of adding position embeddings to the input, ALiBi adds a fixed set of static, non-learned biases to each attention calculation to incorporate positional information into attention scores. Specifically, let $\mathbf{q}_{i}$ define the $i$-thth query in the input sequence of length $L$ and $\mathbf{K}$ defines the key matrix, the attention score of query $i$ is calculated as: $\texttt{softmax}(\mathbf{q}_{i}\mathbf{K}+m*[-(i-1),...,-2,-1,0,-1,-2,...,-(L-1-i)])$, where $m$ is a fixed head-specific constant. ALiBi used a geometric sequence (i.e., $\frac{1}{2^{1}},\frac{1}{2^{2}},...,\frac{1}{2^{n}}$) of different $m$ to each attention head. Intuitively, ALiBi increasingly penalizes attention scores between key-query pairs as their distances increase, and $m$ determines the penalty rate. Replacing learned position embedding with ALiBi allows DNABERT-2 to effectively handle arbitrarily long sequences during fine-tuning and inference despite being pre-trained on relatively short sequences.
@@
Flash attention is an IO-aware algorithm that implements the exact standard attention calculation in a more time- and memory-efficient way. It identifies a main bottleneck of standard attention implementation as the lack of taking the number of reads and writes to fast GPU on-chip SRAM and relatively slow GPU high bandwidth memory (HBM) into account. To avoid reading and writing to the slow HBM, it splits Key/Query/Value matrices into blocks and incrementally performs softmax over the entire input. It also proposes to recompute large intermediate results like attention scores in backward pass to trade extra computation for fewer IO with HBM, which empirically leads to less computational time. It accelerates DNABERT-2 without sacrificing model performance.
@@
Besides, we replace the Relu activation function with GEGLU [Shazeer, 2020], a variant of GLU [Dauphin et al., 2017] that has been shown to improve the performance of Transformer models. The GEGLU function is calculated as $\texttt{GEGLU}(x,W,V,b,c)=\texttt{GELU}(xW+b)\otimes(xV+c)$, where $x$ is the function input, $W$ and $V$ are learnable weights, and $b$ and $c$ are learnable biases. The GELU function is defined as $\texttt{GELU}(x)=x\Phi(x)$, where $\Phi(x)$ is the cumulative distribution function (CDF) of the standard normal distribution.
3.3 Implementation
We pre-train DNABERT-2 with the Masked Language Modeling (MLM) loss with a mask ratio of $15\%$. Notably, we independently mask every token instead of masking spans of continuous tokens like Ji et al. [2021]. We use a batch size of $4096$ and a max sequence length of $128$. We train the model for $500000$ steps using the AdamW [Loshchilov and Hutter, 2019] optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.98$, $\epsilon=1e\texttt{-}6$ and weight decay of $1e\texttt{-}5$. The learning rate linearly increases from 0 to $5e\texttt{-}4$ during the first $30000$ steps while linearly decreasing to 0 in the last $470000$ steps. The pre-training stage takes approximately 14 days using eight Nvidia RTX 2080Ti GPUs. To train the model, we used the Transformers library by HuggingFace [Wolf et al., 2020] and the Composer library by MosaicML [Team, 2021].
@@ 4.1 Pre-Train: Human and Multi-Species Genome
To investigate the impact of species diversity on genome foundational models, we’ve compiled and made publicly available two datasets for foundational model pre-training: the human genome and the multi-species genome. The human genome dataset is borrowed from DNABERT’s pre-training data [Ji et al., 2021], which comprises 2.752.75B nucleotide bases. The multi-species genome dataset encompasses genomes from $135$ species, spread across $7$ categories. In total, this dataset includes 32.4932.49B nucleotide bases, nearly $12$ times the volume of the human genome dataset. We exclude all sequences with N and retain only sequences that consist of A, T, C, and G. Detailed statistics are presented in Table LABEL:tb:multi_species_details.
4.2 Benckmark: Genome Understanding Evaluation (GUE)
@@
We introduce the Genome Understanding Evaluation (GUE) benchmark, which includes $7$ genome sequence classification problems with $28$ datasets. Table 1 presents the summarization of the GUE benchmark. To evaluate models’ capability in modeling sequences with different lengths, we select datasets with input lengths ranging from $70$ to $1000$. GUE contains tasks for $4$ species: human, mouse, virus, and yeast, to evaluate the multi-species transferability in genome understanding of each model. We explicitly define evaluation metrics for each task and split each dataset into training, validation, and test data for a fair comparison across different models.
To calibrate the GUE benchmark’s difficulty level and better illuminate each model’s capabilities, we carefully selected datasets that are neither too simple nor overly challenging for current models. For example, when the Nucleotide Transformer variants [Lopez et al., 2023] were tested on the Splice Site Prediction dataset, all variants achieved an accuracy between $97\%$ and $98\%$. Similar outcomes were observed in tasks like Promoter Prediction and Enhancer Prediction. These high scores might suggest these variants perform similarly, but as our experiments in Section 5 show, they vary significantly on more discerning datasets.
The construction of GUE starts with the aggregation of various biologically important genome analysis datasets, followed by the assessment of existing models such as DNABERT [Ji et al., 2021] and Nucleotide Transformer [Lopez et al., 2023] on these datasets. Datasets where the majority of models yielded moderate (e.g., F1-scores between 0.3 and 0.8) and distinguishable performance scores were retained. On the other hand, datasets that did not meet these criteria underwent a restructuring process involving various strategies such as class balancing, adversarial sample inclusion, and reduction of training sample volume, among others. After several iterations of this process, we ultimately arrived at 28 representative datasets of moderate difficulty. Due to space limits, we present the detailed data processing and statistics of each dataset in Sec. LABEL:subsec:appen_data_gue.
@@ 5 Experiments
We evaluate DNABERT-2 using the proposed GUE benchmark to thoroughly investigate its versatility and robustness across a variety of tasks involving multi-species genomes.
@@ 5.1 Baseline
We compare DNABERT-2 with two state-of-the-art genome foundation models: DNABERT [Ji et al., 2021] and Nucleotide Transformer [Lopez et al., 2023].
@@
Nucleotide Transformer (NT) scales up the data and model size to achieve state-of-the-art performance in $27$ DNA analysis tasks. It also has $4$ variants: NT-500M-human, NT-500M-1000g, NT-2500M-1000g, and NT-2500M-multi, where human, 1000g, and multi respectively refers to the GRCh38/hg38 human reference genome, 3202 high-coverage human genomes from the 1000 Genome project [Byrska-Bishop et al., 2021], and genome from 850 different species.
It is important to note that NT models are 6 to 29 times larger than DNABERT, which precludes standard model fine-tuning on consumer GPUs. Therefore, we perform standard fine-tuning for DNABERT and DNABERT-2, while adapting the Low-Rank Adaptation (LoRA) technique for fine-tuning the Nucleotide Transformer to enhance efficiency. For a fair comparison, we conducted preliminary experiments to confirm that our implementation of NT achieves comparable results to those reported in their original paper [Lopez et al., 2023] (see Appendix A.3 for more details).
@@ 5.2 Setup and Metric
We evaluate the models from two perspectives: computational efficiency and performance on downstream tasks. To measure each model’s computational cost, we consider the number of model parameters and the relative Floating Point Operations (FLOPs)—which is the total number of multiplication and addition operations during a forward pass—compared to DNABERT-2. We evaluate FLOPs on genome sequences with a length of 500, a commonly used setup in genome analysis. To measure model performance, we utilize F1-Score and Matthews Correlation Coefficient (MCC). We use different metrics for different tasks, following conventional practices (refer to Table LABEL:tb:gue_details for details). Table 2 presents the overall performance of each model on the GUE benchmark. It provides the average score of each model and the number of times it ranks in the top two among all models. The average results across all tasks are reported in Table 3, while task-specific results can be found in 4, with full details relegated to this section due to space constraints. We also include statistics on the number of tokens each model processed during its pre-training phase, providing insight into the effects of training steps on model performance. For each model, we keep most of the hyperparameters (e.g., learning rate, batch size, weight decay, etc.) constant across all datasets, adjusting only the maximum sequence length and the number of training steps according to the specific dataset. Hyperparameter tuning tailored to each dataset is left for future work. Throughout the training process, we validate the model every 200 steps, save the model that yields the smallest loss on the validation set, and report its evaluation results on the test set. We train each model using three different random seeds and report the average results.
@@
We also investigate the impact of additional in-domain pre-training on DNA language models. We combine the training sets of the 28 GUE datasets and further pre-train DNABERT-2 on this combined set. Following Sun et al. [2020], we train the model with a batch size of 32, a maximum sequence length of 128, and a learning rate of $5e\texttt{-}5$ for 100,000 steps. This results in the processing of 0.41B tokens, which only constitute 0.08% of the tokens processed during the entire training process of DNABERT-2.
5.3 Main Results
Model Num. Params. $\downarrow\downarrow$ FLOPs $\downarrow\downarrow$ Trn. Tokens Num. Top-2 $\uparrow\uparrow$ Ave. Scores $\uparrow\uparrow$
@@ 5.3 Main Results
Yeast Mouse Virus Human
EMP TF-M CVC TF-H PD CPD SSP
@@
Table 3: The models’ averaged performance on the 88 tasks in the GUE benchmark, including Epigenetic Marks Prediction (EMP), Transcription Factor Prediction on the Human genome and the Mouse genome (TF-H and TF-M), Covid Variants Classification (CVC), Promoter Detection (PD), Core Promoter Detection (CPD), and Splice Site Prediction (SSP).
Although DNABERT-2 is $5$ times smaller, it surpasses NT-500M while using less FLOPs. This underscores the importance of providing the model with adequate data, particularly when the model size is scaled up, and further highlights the inefficiency of overlapping k-mer tokenization. The comparison between DNABERT and NT-2500M-1000g exposes the sample inefficiency of non-overlapping k-mer tokenization. Despite being trained on $2.5$ times more tokens, NT-2500M-1000g achieves a performance similar to that of DNABERT.
The averaged results for each task are displayed in Table 3. DNABERT-2 and NT-2500M-multi consistently achieve top-tier performance across most tasks. Their dominance over other baselines is particularly notable in non-human genome analysis tasks, demonstrating the effectiveness of pre-training on multi-species genomes. Furthermore, models trained on multi-species genomes also show strong performance on human genome analysis tasks, proving their ability to develop a comprehensive understanding of multi-species genomes without compromising their grasp of the human genome. However, we observe that additional pre-training does not uniformly enhance performance across all tasks, indicating that task-specific further pre-training might be beneficial when addressing a certain downstream task.
@@ 6 Conclusion
In this paper, we introduce DNABERT-2, an efficient foundational model pre-trained on extensive multi-species genomes. We identify the computational and sample inefficiencies of the existing k-mer tokenization method and propose the adaptation of Byte Pair Encoding (BPE) for DNA language modeling. We provide insightful and comprehensive empirical analyses, building DNABERT-2 based on these findings. Moreover, we integrate several techniques such as Attention with Linear Biases (ALiBi) and Low-Rank Adaptation (LoRA) to address the limitations of current DNA language models. From a data perspective, we compile and introduce the Genome Understanding Evaluation (GUE), a benchmark for multi-species genome analysis comprising seven tasks and 28 datasets with well-defined training, validation, test sets, clear evaluation metrics, and elaborately calibrate difficulty. In addition, we release a multi-species genome dataset consisting of $32.49$ billion nucleotide bases derived from the genomes of 135 species across seven categories. We believe these datasets will significantly contribute to the progression of research on DNA language models. For future work, we identify several promising directions: 1) the development of effective modeling strategies for short genome sequences; 2) scaling up the model size; and 3) the introduction of training targets and data processing/augmentation methods that leverage the unique double-strand structure of DNA.
Acknowledgments

### 18 Mar 2024 v2

@@ 3.2 Model
DNABERT-2 adapts the Transformer Encoder architecture. To address the limitations of existing models, we incorporate a series of recent advances in deep learning to increase the model’s efficiency and capability, including: 1) replacing learned positional embeddings with the Attention with Linear Biases (ALiBi) (Press et al., 2021) to overcome the input length limitation; 2) utilizing FlashAttention (Dao et al., 2022) and Low Precision Layer Normalization to increase computation and memory efficiency; 3) employing the Low-Rank Adaptation (LoRA) (Hu et al., 2021) in the fine-tuning stage (if necessary) for parameter-efficient training.
@@
Due to the permutation-invariant nature of the attention mechanism, explicit positional information is required in attention-based models. Existing solutions such as Sinusoidal (Vaswani et al., 2017), learned (Kenton & Toutanova, 2019), and Rotary (Su et al., 2021) positional embedding methods either suffer from input length restriction or poor extrapolation capability when applied to sequences longer than training data. Attention with Linear Biases (ALiBi) provides an efficient yet effective solution. Instead of adding position embeddings to the input, ALiBi adds a fixed set of static, non-learned biases to each attention calculation to incorporate positional information into attention scores. Specifically, let $\mathbf{q}_{i}$ define the $i$-thth query in the input sequence of length $L$ and $\mathbf{K}$ defines the key matrix, the attention score of query $i$ is calculated as: $\texttt{softmax}(\mathbf{q}_{i}\mathbf{K}+m*[-(i-1),...,-2,-1,0,-1,-2,...,-(L-1-i)])$, where $m$ is a fixed head-specific constant. ALiBi used a geometric sequence (i.e., $\frac{1}{2^{1}},\frac{1}{2^{2}},...,\frac{1}{2^{n}}$) of different $m$ to each attention head. Intuitively, ALiBi increasingly penalizes attention scores between key-query pairs as their distances increase, and $m$ determines the penalty rate. Replacing learned position embedding with ALiBi allows DNABERT-2 to effectively handle arbitrarily long sequences during fine-tuning and inference despite being pre-trained on relatively short sequences.
@@
Flash attention is an IO-aware algorithm that implements the exact standard attention calculation in a more time- and memory-efficient way. It identifies a main bottleneck of standard attention implementation as the lack of taking the number of reads and writes to fast GPU on-chip SRAM and relatively slow GPU high bandwidth memory (HBM) into account. To avoid reading and writing to the slow HBM, it splits Key/Query/Value matrices into blocks and incrementally performs softmax over the entire input. It also proposes to recompute large intermediate results like attention scores in backward pass to trade extra computation for fewer IO with HBM, which empirically leads to less computational time without sacrificing model performance.
@@
Besides, we replace the Relu activation function with GEGLU (Shazeer, 2020), a variant of GLU (Dauphin et al., 2017) that has been shown to improve the performance of Transformer models. The GEGLU function is calculated as $\texttt{GEGLU}(x,W,V,b,c)=\texttt{GELU}(xW+b)\otimes(xV+c)$, where $x$ is the function input, $W$ and $V$ are learnable weights, and $b$ and $c$ are learnable biases. The GELU function is defined as $\texttt{GELU}(x)=x\Phi(x)$, where $\Phi(x)$ is the cumulative distribution function (CDF) of the standard normal distribution.
@@ 4.1 Pre-Train: Human and Multi-Species Genome
To investigate the impact of species diversity on genome foundation models, we compile and made publicly available two datasets for foundation model pre-training: the human genome and the multi-species genome. The human genome dataset is the one used in DNABERT (Ji et al., 2021), which comprises 2.752.75B nucleotide bases. The multi-species genome dataset encompasses genomes from $135$ species, spread across $6$ categories. In total, this dataset includes 32.4932.49B nucleotide bases, nearly $12$ times the volume of the human genome dataset. We exclude all sequences with N and retain only sequences that consist of A, T, C, and G. Detailed statistics are presented in Table LABEL:tb:multi_species_details.
4.2 Benckmark: Genome Understanding Evaluation (GUE & GUE+)
@@
Species Task Num. Datasets Num. Classes Sequence Length
Human Enhancer Promoter Interaction 6 2 5000
Fungi Species Classification 1 25 5000
Virus Species Classification 1 20 10000
Table 2: Summarization of the Genome Understanding Evaluation Plus (GUE+) benchmark.
We introduce a large benchmark for genome foundation models. Due to the input length limits of existing genome foundation models, we split the benchmark into two parts: GUE and GUE+. GUE (See Table 1) includes $7$ genome sequence classification problems with $28$ datasets with input lengths ranging from $70$ to $1000$. GUE+ (See Table 2), on the other hand, focuses on relatively longer input sequences, rating from $5000$ to $10000$. To evaluate the multi-species transferability in genome understanding of each model, this benchmarks contains tasks for various species, including human, fungi, virus, and yeast. We explicitly define evaluation metrics for each task and split each dataset into training, validation, and test data for a fair comparison across different models.
To calibrate the benchmark’s difficulty level and better illuminate each model’s capabilities, we carefully selected datasets that are neither too simple nor overly challenging for current models. For example, when the Nucleotide Transformer variants (Dalla-Torre et al., 2023) were tested on the Splice Site Prediction dataset, all variants achieved an accuracy between $97\%$ and $98\%$. Similar outcomes were observed in tasks like Promoter Prediction and Enhancer Prediction. These high scores might suggest these variants perform similarly, but as our experiments in Section 5 show, they vary significantly on more discerning datasets.
The construction of the benchmark starts with the aggregation of various biologically important genome analysis datasets, followed by the assessment of existing models such as DNABERT (Ji et al., 2021) and Nucleotide Transformer (Dalla-Torre et al., 2023) on these datasets. Datasets where the majority of models yielded moderate (e.g., F1-scores between 0.3 and 0.8) and distinguishable performance scores were retained. On the other hand, datasets that did not meet these criteria underwent a restructuring process involving various strategies such as class balancing, adversarial sample inclusion, and reduction of training sample volume, among others. After several iterations of this process, we ultimately arrived at 36 representative datasets of moderate difficulty. Due to space limits, we present the detailed data processing and statistics of each dataset in Sec. LABEL:subsec:appen_data_gue.
@@ 5 Experiments
We evaluate DNABERT-2 using the proposed benchmark to thoroughly investigate its versatility and robustness across a variety of tasks involving multi-species genomes.
@@ 5.1 Baseline
We compare DNABERT-2 with two state-of-the-art genome foundation models: DNABERT (Ji et al., 2021) and Nucleotide Transformer (Dalla-Torre et al., 2023).
@@
Nucleotide Transformer (NT) scales up the data and model size to achieve state-of-the-art performance in $27$ DNA analysis tasks. It also has $4$ variants: NT-500M-human, NT-500M-1000g, NT-2500M-1000g, and NT-2500M-multi, where human, 1000g, and multi respectively refers to the GRCh38/hg38 human reference genome, 3202 high-coverage human genomes from the 1000 Genome project (Byrska-Bishop et al., 2021), and genome from 850 different species.
It is important to note that NT models are 6 to 29 times larger than DNABERT, which precludes standard model fine-tuning on consumer GPUs. Therefore, we perform standard fine-tuning for DNABERT and DNABERT-2, while adapting the Low-Rank Adaptation (LoRA) technique for fine-tuning the Nucleotide Transformer to enhance efficiency. For a fair comparison, we conducted preliminary experiments to confirm that our implementation of NT achieves comparable results to those reported in their original paper (Dalla-Torre et al., 2023) (see Appendix A.4 for more details).
@@ 5.2 Setup and Metric
We evaluate the models from two perspectives: computational efficiency and performance on downstream tasks. To measure each model’s computational cost, we consider the number of model parameters and the relative Floating Point Operations (FLOPs)—which is the total number of multiplication and addition operations during a forward pass—compared to DNABERT-2. We evaluate FLOPs on genome sequences with a length of 500, a commonly used setup in genome analysis. To measure model performance, we utilize F1-Score and Matthews Correlation Coefficient (MCC). We use different metrics for different tasks, following conventional practices (refer to Table LABEL:tb:gue_details for details). Table 3 presents the overall performance of each model on the GUE benchmark. It provides the average score of each model and the number of times it ranks in the top two among all models. The average results across all tasks are reported in Table 4, while task-specific results are presented in Table 6 due to space constraints. We also include statistics on the number of tokens each model processed during its pre-training phase, providing insight into the effects of training steps on model performance. For each model, we keep most of the hyperparameters (e.g., learning rate, batch size, weight decay, etc.) constant across all datasets, adjusting only the maximum sequence length and the number of training steps according to the specific dataset. Hyperparameter tuning tailored to each dataset is left for future work. Throughout the training process, we validate the model every 200 steps, save the model that yields the smallest loss on the validation set, and report its evaluation results on the test set. We train each model using three different random seeds and report the average results.
@@
We also investigate the impact of additional in-domain pre-training on DNA language models. We combine the training sets of the 28 GUE datasets and further pre-train DNABERT-2 on this combined set. Following Sun et al. (2020), we train the model with a batch size of 32, a maximum sequence length of 128, and a learning rate of $5e\texttt{-}5$ for 100,000 steps. This results in 0.41B training tokens, which only constitute 0.08% of the tokens processed during the entire training process of DNABERT-2.
5.3 Results on GUE
Model Params. $\downarrow\downarrow$ FLOPs $\downarrow\downarrow$ Trn. Tokens Num. Top-2 $\uparrow\uparrow$ Ave. Scores $\uparrow\uparrow$
@@ 5.3 Results on GUE
Although DNABERT-2 is $5$ times smaller, it surpasses NT-500M while using fewer FLOPs. This underscores the importance of providing the model with adequate data, particularly when the model size is scaled up, and further highlights the inefficiency of overlapping k-mer tokenization. The comparison between DNABERT and NT-2500M-1000g exposes the sample inefficiency of non-overlapping k-mer tokenization. Despite being trained on $2.5$ times more tokens, NT-2500M-1000g achieves a performance similar to that of DNABERT.
Species Yeast Mouse Virus Human
Task EMP TF-M CVC TF-H PD CPD SSP
@@
Table 4: The models’ averaged performance on the 77 tasks in the GUE benchmark, including Epigenetic Marks Prediction (EMP), Transcription Factor Prediction on the Human genome and the Mouse genome (TF-H and TF-M), Covid Variants Classification (CVC), Promoter Detection (PD), Core Promoter Detection (CPD), and Splice Site Prediction (SSP).
The averaged results for each task are displayed in Table 4. DNABERT-2 and NT-2500M-multi consistently achieve top-tier performance across most tasks. Their dominance over other baselines is particularly notable in non-human genome analysis tasks, demonstrating the effectiveness of multi-species genomes pre-training. Furthermore, models trained on multi-species genomes also show strong performance on human genome analysis tasks, proving their ability to develop a comprehensive understanding of multi-species genomes without compromising their grasp of the human genome. We also observe that additional pre-training does not uniformly enhance all tasks, indicating that task-specific further pre-training might be necessary when addressing a certain downstream task.
@@
5.4 Results on GUE+
Task SC (Fungi) SC (Virus) EPI (Human)
Dataset 0 1 2 3 4 5
DNABERT (6-mer) 89.29 44.51 - - - - - -
NT-2500M-multi 92.85 45.00 61.91 72.15 73.13 79.49 86.48 68.64
DNABERT-2 93.04 48.50 76.21 79.19 83.50 86.71 92.90 73.70
Table 5: The models’ averaged performance on the 33 tasks in the GUE+ benchmark, including 66 Enhance Promoter Interaction (EPI) datasets and two Species Classification (SC) datasets on virus and fungi. The - means we are not able to fit the model on the dataset after trying multiple sets of hyperparameters.
In this section, we further compare DNABERT-2 with DNABERT and Nucleotide Transformers on the GUE+ benchmark. The input size of datasets in GUE+ ranges from $5000$ to $10000$. For DNABERT, due to its 512 input limitation, we follow its origin setting by splitting the input into 512bp-len pieces, generating embedding independently, and feeding the averaged embedding to the output layer. Table 5 shows DNABERT, Nucleotide Transformer, and DNABERT-2’s performance on the GUE+ benchmark. As shown in the table, DNABERT-2 consistently outperforms the baselines on all the datasets, showing its good capability in handling long DNA sequences. Notably, despite DNABERT-2 is pre-trained purely on 700bp-len sequences, it performs well even on 10000-bp sequences with a few epochs of fine-tuning, indicating DNABERT-2’s extrapolation capability offered by ALiBi.
@@ 6 Conclusion
In this paper, we introduce DNABERT-2, an efficient foundational model pre-trained on multi-species genomes. We identify the computational and sample inefficiencies of the existing k-mer tokenization method and propose the adaptation of Byte Pair Encoding (BPE) for DNA language modeling. We provide insightful and comprehensive empirical analyses, building DNABERT-2 based on these findings. Moreover, we integrate several techniques such as Attention with Linear Biases (ALiBi) and Low-Rank Adaptation (LoRA) to address the limitations of current DNA language models. We also compile and introduce the Genome Understanding Evaluation, a benchmark for multi-species genome analysis comprising $9$ tasks and $36$ datasets with well-defined data split, evaluation metrics, and elaborately calibrated difficulty. Empirical analysis on the GUE benchmark shows that DNABERT-2 achieves comparable performance as the SOTA models while being much more efficient. For future work, we are interested in effective modeling strategies for short and extra-long genome sequences and the introduction of training targets and data processing/augmentation methods that leverage the unique double-strand structure of DNA.
Acknowledge